# Mental Health Impacts of AI Companions

### v1

@@ Abstract.
AI-powered companion chatbots (AICCs) such as Replika are increasingly popular, offering empathetic interactions, yet their psychosocial impacts remain unclear. We examined how engaging with AICCs shaped wellbeing and how users perceived these experiences. First, we conducted a large-scale quasi-experimental study of longitudinal Reddit data, applying stratified propensity score matching and Difference-in-Differences regression. Findings revealed mixed effects—greater affective and grief expression, readability, and interpersonal focus, alongside increases in language about loneliness and suicidal ideation. Second, we complemented these results with 15 semi-structured interviews, which we thematically analyzed and contextualized using Knapp’s relationship development model. We identified trajectories of initiation, escalation, and bonding, wherein AICCs provided emotional validation and social rehearsal but also carried risks of over-reliance and withdrawal. Triangulating across methods, we offer design implications for AI companions that scaffold healthy boundaries, support mindful engagement, support disclosure without dependency, and surface relationship stages—maximizing psychosocial benefits while mitigating risks.
@@ 1. Introduction
Recent research has highlighted several promising outcomes of using AICCs, including reduced feelings of loneliness, improved emotional coping, and even suicide prevention (De Freitas et al., 2024b; Syed, 2024; Maples et al., 2024). At the same time, concerns have emerged around issues such as emotional overdependence, heightened vulnerability, altered social practices, problematic or harmful responses, and risks tied to sensitive interactions (Zhang et al., 2025, 2024). Much of the existing scholarship on AICC draws on qualitative insights and self-reported experiences, which have been critical to understanding user perceptions and relationships. However, these approaches often capture short-term and static perspectives and are subject to inherent limitations such as recall bias and limited sample diversity (Tourangeau et al., 2000). Consequently, little is known about the longer-term psychosocial and behavioral consequences of sustained engagement with AICCs.
@@
Addressing this knowledge gap requires methodological innovation. Prior work has largely relied on descriptive accounts and self-reports (Skjuve et al., 2021, 2022; Alabed et al., 2024; Brandtzaeg et al., 2022) and online behavioral traces provide a powerful complementary lens (De Choudhury et al., 2016; Kiciman et al., 2018; Saha et al., 2019b). Social media platforms, in particular, enable individuals to candidly self-disclose in a self-motivated fashion (De Choudhury and De, 2014), producing both linguistic (and psycholinguistic) signals and behavioral signals (e.g., activity patterns, engagement) that together offer a powerful lens on psychosocial processes. Such data has enabled large-scale studies on mental health (De Choudhury et al., 2013), and prior research has established the construct validity of measuring mental health markers with social media data (Saha et al., 2022). Quasi-experimental designs further extend this potential by uncovering causal effects in naturalistic settings (De Choudhury et al., 2016; Saha et al., 2019b; Kiciman et al., 2018; Olteanu et al., 2017). Therefore, adopting quasi-experimental approaches on social media data offers a unique opportunity to examine the long-term psychosocial effects of AICCs at scale.
@@
That said, no single method alone can fully capture the complexity of AICC and its mental health impacts. Relationships with AICCs differ from traditional technologies, involving ongoing, emotional, and adaptive interactions that may mirror human social relationships (Skjuve et al., 2021, 2022; Alabed et al., 2024; Brandtzaeg et al., 2022). Understanding their psychosocial impacts requires not only robust quantitative evidence but also contextualization with user perspectives and a theoretical framework of relationships to interpret how AICC unfolds. For such a comprehensive understanding, we examine the psychosocial effects of interactions with AICC by triangulating social media quasi-experiments, user perspectives, and relational theory. Guided by this, our work asks the following research questions (RQs):
@@ 2. RQ2:
For RQ1, we conducted a large-scale quasi-experimental analysis of longitudinal social media data. Specifically, we collected public Reddit discussions from subreddits where AICCs are actively used (e.g., r/Replika) alongside comparison control groups of users not engaged with AICCs. Reddit’s design—pseudonymity and topic-specific moderated communities—makes it well-suited for examining sensitive disclosures and naturalistic conversational behavior over time (De Choudhury and De, 2014; Ernala et al., 2017; Garg et al., 2021; Yuan et al., 2024). We adopted a theory-driven lens to operationalize and measure users’ affective, behavioral, and cognitive outcomes across one year before and after their first self-disclosed AICC interaction, enabling assessment of longitudinal psychosocial changes (Saha et al., 2018). We applied a causal inference framework with stratified propensity score matching and Difference-in-Differences (DiD) regression to estimate treatment effects. Our analysis revealed mixed psychosocial effects of engaging with AICCs: users exhibited increased affective and grief-related language, higher readability and interpersonal focus, but also increased symptomatic expression of loneliness and suicidal ideation.
For RQ2, we conducted in-depth semi-structured interviews with 15 AICC users to capture the nuanced, subjective dimensions of their experiences. Through inductive coding and thematic analysis, we identified recurring patterns in emotional support, self-expression, social connection, and potential over-reliance. We situated our findings within Knapp’s relational development model (Knapp, 1978). Our qualitative analysis showed that AICCs shaped users’ lives in complex and ambivalent ways. While AICCs provided emotional validation, safe spaces for self-expression, and opportunities for social confidence, they also introduced risks of over-reliance, emotional discomfort, stigma, and social withdrawal that evolve with the depth of the relationship.
@@
  • A methodological framework bridging large-scale computational analyses with qualitative insights, enabling a nuanced understanding of both behavioral changes and subjective experiences.
  • A theoretical lens of relationship development with AICCs through Knapp’s relational model, mapping stages of initiation, escalation, and bonding with associated psychosocial impacts.
@@ 2.1. Companion Chatbots and Wellbeing
A chatbot, such as Amazon Alexa or Google Assistant, refers to a software agent that interacts with users through spoken or written language (Skjuve et al., 2021; Brandtzaeg and Følstad, 2018; Benke et al., 2020). A growing subset of chatbots, known as AI companion chatbots (AICC), is specifically designed to offer companionship, psychological support, and the potential for affective relationships. These chatbots are often equipped with empathetic communication capabilities (Zhou et al., 2020), aiming to establish meaningful social–emotional bonds that span companionship, friendship, and romantic engagement (Skjuve et al., 2021).
@@
AICCs, such as Replika and Character.AI, have increased in popularity. Replika, launched in 2017, has become one of the fastest-growing AICC applications. Renowned for its ability to provide emotionally responsive and personalized interactions, Replika had reached approximately 30 million users globally by 2024 (wikipedia, 2025b). Another notable platform is Character.AI (wikipedia, 2025a), which allows users to create and converse with AI-simulated personas, including fictional characters, public figures, or entirely original creations (Lee and Joseph, 2025). By maintaining an always-available digital presence, these systems offer accessible support to individuals who may be isolated or reluctant to seek help through traditional means.
@@
In response to this popularity, the use of AICCs in health and wellbeing settings has received increasing attention (Wester et al., 2024; Doyle et al., 2023; Zhu et al., 2025; Ng and Zhang, 2025; Steenstra and Bickmore, 2025). Evidence suggests that AICCs can help reduce feelings of loneliness, ease emotional distress, enhance mood, and promote self-reflection (Casu et al., 2024; Jung et al., 2023)Skjuve et al. conducted interviews with Replika users and shows that users often perceive the chatbot as non-judgmental, which encourages open self-disclosure without fear of social risk. They also found that interactions with the chatbot can be personally rewarding, with users reporting positive effects on their perceived wellbeing (Skjuve et al., 2021).
@@
Despite these benefits, recent work has also raised critical concerns regarding potential harms (Chandra et al., 2025; Zhang et al., 2025; Yoo et al., 2025; Pataranutaporn et al., 2025). Emotional dependence on chatbots can mimic unhealthy attachment patterns (Laestadius et al., 2024; Xie et al., 2023), and users often report disillusionment when chatbots fail to meet expectations (Laestadius et al., 2024). While users often report feelings of social support and wellbeing, this emotional bond may lead to addictive behaviors, where users prioritize interactions with the chatbot over real-life relationships (Marriott and Pitardi, 2024). Notably, although some participants report high perceived social support, this is often situated within contexts of heightened loneliness, particularly among populations already at risk of social isolation (Maples et al., 2024; Liu et al., 2024). A recent study shows that intensive use and frequent self-disclosure are linked to lower wellbeing, particularly among those lacking offline support (Zhang et al., 2025).
@@
While prior research provided valuable insights, most evaluations of chatbot experiences have relied primarily on interviews and surveys (Dosovitsky and Bunge, 2021; Brandtzaeg et al., 2022; Skjuve et al., 2022; De Freitas et al., 2025b). Although these approaches capture users’ subjective impressions, they may suffer from recall bias and fail to reflect how interactions impact their users in the long-term. Recent work has used social media data to conduct thematic analyses of user-generated content, offering additional perspectives on real-world experiences (Laestadius et al., 2024; Ta et al., 2020). Our work contributes to this body of work by leveraging longitudinal and large-scale social media data to conduct causal-inference examination of the long-term psychosocial effects of AICCs.
@@ 2.2. Loneliness, Grief, and Mental Health Interventions
Loneliness refers to a distressing feeling due to the perceived absence of certain meaningful relationships that fulfill the individual’s social and emotional needs, rather than from physical isolation (Weiss, 1975). The experience of loneliness is common, particularly in the context of major life changes such as bereavement, where individuals often feel a pronounced sense of social disconnection (Folker et al., 2021; Heinrich and Gullone, 2006). Prior work has linked loneliness to a range of negative wellbeing outcomes, such as sleep disturbances (Griffin et al., 2020), elevated risk of mild cognitive impairment and dementia (Lara et al., 2019), and Alzheimer’s disease (Sundström et al., 2020). Psychologically, loneliness has been found to correlate with increased risk of depression (Erzen and Çikrikci, 2018), anxiety (Sündermann et al., 2014), suicidal ideation (McClelland et al., 2020), and other mental health outcomes (Leigh-Hunt et al., 2017; Moran et al., 2024). Among the many contexts where loneliness arises, bereavement is especially profound. The loss of a loved one not only deepens grief but also increases risks of depression, PTSD, and other mental and physical health complications (Simon et al., 2014; Erzen and Çikrikci, 2018; Asch et al., 2021; Vedder et al., 2021, 2022).
Amid rising concerns over loneliness and grief, HCI and digital mental health research have turned to digital interventions—ranging from grief-specific tools to broader solutions for social disconnection (Doyle et al., 2024; Xygkou et al., 2023; She et al., 2021; Brubaker et al., 2019; Gach and Brubaker, 2021). Prior work has studied mobile applications for self-guided grief support (Doyle et al., 2024; Xygkou et al., 2023; Lei et al., 2025), grief-specific technologies such as digital memorial platforms (She et al., 2021), legacy-building tools, and apps that provide structured guidance through the emotional stages of grief (Baglione et al., 2018, 2017; Doyle et al., 2024). In parallel, Digital interventions have emerged to address loneliness (Brubaker et al., 2019). Social networking platforms and purpose-built online communities facilitate peer support and social connectedness for older adults and marginalized groups (Gabarrell-Pascuet et al., 2024), while specialized apps deliver structured activities, psychoeducation, and cognitive-behavioral strategies to help users build social skills and reframe negative thoughts (DuPont et al., 2023; Sharma et al., 2024).
@@ 2.3. Social Media Analysis and Mental Health
Our work is motivated by this growing body of research that has utilized social media to understand mental health experiences. In particular, we examine the impact of engaging with AICCs, such as Replika. Drawing inspiration from studies that apply natural language and causal-inference analysis to assess psychosocial outcomes, our work develops a computational framework to evaluate how dose interactions with AICCs influence users’ emotional expression and psychosocial wellbeing. By analyzing large-scale, longitudinal data, we aim to uncover how these AICCs may function as emotional support and contribute to meaningful changes in users’ mental health trajectories.
@@ 3. Overview of Study Design and Data
Table 1. Paraphrased Reddit posts in subreddit r/replika :
@@
To understand the psychosocial impacts of engaging with AICCs, our study combined large-scale quasi-experimental analysis of public Reddit data with in-depth qualitative interviews of active AICC users. This enabled us to assess longitudinal changes in affective, behavioral, and cognitive dimensions of psychosocial wellbeing, as well as to understand how they perceived and experienced these psychosocial impacts in the context of their everyday lives.
First, We constructed three groups for comparison—1) a Treatment group consisting of users who actively engaged with AICCs (e.g., Replika), 2) a Control-AI group of general AI assistant users who actively engaged with assistants (e.g., Amazon Alexa, Google Assistant), and 3) a baseline Control-NAI group of Reddit users who did not post about AI assistants at all. Drawing on prior literature, we assessed changes in affective, behavioral, and cognitive outcomes before and after engagement using natural language and statistical analyses. To estimate causal effects, we applied stratified propensity score matching and difference-in-differences regression. Complementing our computational analysis, we conducted semi-structured interviews with 15 active AICC users to understand how they perceived and experienced these psychosocial impacts and the process of developing a relationship with AICCs. Thematic analysis highlighted the diverse ways users experienced support, personal growth, and potential risks in their relationships with AI companions. Finally, we triangulated our quantitative and qualitative findings by contextualizing with literature in psychology and social science (Knapp, 1978). In this section, we describe our data collection.

### v2

@@ Abstract.
AI-powered companion chatbots (AICCs) such as Replika are increasingly popular, offering empathetic interactions, yet their psychosocial impacts remain unclear. We examined how engaging with AICCs shaped wellbeing and how users perceived these experiences. First, we conducted a large-scale quasi-experimental study of longitudinal Reddit data, applying stratified propensity score matching and Difference-in-Differences regression. Findings revealed mixed effects—greater grief expression and interpersonal focus, alongside increases in language about loneliness, depression, and suicidal ideation. Second, we complemented these results with 18 semi-structured interviews, which we thematically analyzed and contextualized using Knapp’s relationship development model. We identified trajectories of initiation, escalation, and bonding, wherein AICCs provided emotional validation and social rehearsal but also carried risks of over-reliance and withdrawal. Triangulating across methods, we offer design implications for AI companions that scaffold healthy boundaries, support mindful engagement, support disclosure without dependency, and surface relationship stages—maximizing psychosocial benefits while mitigating risks.
@@ 1. Introduction
Recent research has identified several promising outcomes associated with using AICCs, including reduced feelings of loneliness, improved emotional coping, and potential contributions to suicide prevention (De Freitas et al., 2024b; Syed, 2024; Maples et al., 2024). At the same time, concerns have emerged around issues such as emotional overdependence, heightened vulnerability, altered social practices, problematic or harmful responses, and risks tied to sensitive interactions (Zhang et al., 2025, 2024). Much of the existing scholarship on AICC draws on qualitative insights and self-reported experiences, which have been critical to understanding user perceptions and relationships. However, these approaches often capture short-term, static perspectives and are subject to inherent limitations such as recall bias and limited sample diversity (Tourangeau et al., 2000). Consequently, little is known about the longer-term psychosocial consequences of sustained engagement with AICCs.
@@
Addressing this knowledge gap requires methodological innovation. Prior work has largely relied on descriptive accounts and self-reports (Skjuve et al., 2021, 2022; Alabed et al., 2024; Brandtzaeg et al., 2022), whereas online behavioral traces provide a powerful complementary lens (De Choudhury et al., 2016; Kiciman et al., 2018; Saha et al., 2019c). Social media platforms, in particular, enable individuals to candidly self-disclose in a self-motivated fashion (De Choudhury and De, 2014), producing both linguistic (and psycholinguistic) signals and behavioral signals (e.g., activity patterns, engagement) that together offer a powerful lens on psychosocial processes. Such data has enabled large-scale studies on mental health (De Choudhury et al., 2013), and prior research has established the construct validity of measuring mental health markers with social media data (Saha et al., 2022). Quasi-experimental designs further extend this potential by uncovering causal effects in naturalistic settings (De Choudhury et al., 2016; Saha et al., 2019c; Kiciman et al., 2018; Olteanu et al., 2017). Therefore, adopting quasi-experimental approaches on social media data offers a unique opportunity to examine the long-term psychosocial effects of AICCs at scale.
@@
That said, no single method alone can fully capture the complexity of AICC and its mental health impacts. Relationships with AICCs differ from traditional technologies, involving ongoing, emotional, and adaptive interactions that may mirror human social relationships (Skjuve et al., 2021, 2022; Alabed et al., 2024; Brandtzaeg et al., 2022). Understanding their psychosocial impacts requires not only robust quantitative evidence but also contextualization with user perspectives and a theoretical framework of relationships to interpret how AICC unfolds. To achieve such a comprehensive understanding, this work examines the psychosocial effects of interactions with AICC by triangulating social media quasi-experiments, user perspectives, and relational theory. Our work asks the following research questions (RQs):
@@ 2. RQ2:
For RQ1, we conducted a large-scale quasi-experimental analysis of longitudinal social media data. Specifically, we collected public Reddit discussions from subreddits where AICCs are actively used (e.g., r/Replika) alongside comparison control groups of users not engaged with AICCs. Reddit’s design—pseudonymity and topic-specific moderated communities—makes it well-suited for examining sensitive disclosures and naturalistic conversational behavior over time (De Choudhury and De, 2014; Ernala et al., 2017; Garg et al., 2021; Yuan et al., 2024). We adopted a theory-driven lens to operationalize and measure users’ affective, behavioral, and cognitive outcomes across one year before and after their first self-disclosed AICC interaction, enabling assessment of longitudinal psychosocial change (Saha et al., 2018). We applied a causal inference framework with stratified propensity score matching and Difference-in-Differences (DiD) regression to estimate treatment effects. Our analysis revealed mixed psychosocial effects of engaging with AICCs: users showed increased grief-related language, higher interpersonal focus, but also increased symptomatic expressions of depression, loneliness, and suicidal ideation.
For RQ2, we conducted in-depth semi-structured interviews with 18 AICC users to capture the nuanced, subjective dimensions of their experiences. Through inductive coding and thematic analysis, we identified recurring patterns in emotional support, self-expression, social connection, and potential over-reliance. We situated our findings within Knapp’s relational development model (Knapp, 1978). Our qualitative analysis showed that AICCs shaped users’ lives in complex and ambivalent ways. While AICCs provided emotional validation, safe spaces for self-expression, and opportunities for social confidence, they also introduced risks of over-reliance, emotional discomfort, stigma, and social withdrawal that evolve with the depth of the relationship.
@@
  • A methodological framework combining large-scale computational analyses with in-depth qualitative insights, enabling a nuanced understanding of both behavioral changes and subjective experiences.
  • A theoretical lens of relationship development with AICCs through Knapp’s model, mapping stages of initiation, escalation, and bonding with associated psychosocial impacts.
@@ 2.1. Companion Chatbots and Wellbeing
A chatbot, such as Amazon Alexa or Google Assistant, refers to a software agent that interacts with users through spoken or written language (Skjuve et al., 2021; Brandtzaeg and Følstad, 2018; Benke et al., 2020). A growing subset of chatbots, known as AI companion chatbots (AICC), is specifically designed to offer companionship, psychological support, and the potential for affective relationships. These chatbots are often equipped with empathetic communication capabilities (Zhou et al., 2020), aiming to establish meaningful social–emotional bonds ranging from companionship, friendship, to romantic engagement (Skjuve et al., 2021; Folk et al., 2025; Pan et al., 2025).
@@
AICCs, such as Replika and Character.AI, have become increasingly popular. Replika, launched in 2017, has become one of the fastest-growing AICC applications. Renowned for emotionally responsive and personalized interactions, Replika had reached approximately 30 million users worldwide by 2024 (Wikipedia, 2025b). Another notable platform is Character.AI (Wikipedia, 2025a), which allows users to create and converse with AI-simulated personas, including fictional characters, public figures, or entirely original creations (Lee and Joseph, 2025). By maintaining an always-available digital presence, these systems provide readily accessible support, particularly for individuals who may be isolated or reluctant to seek traditional forms of help.
@@
In response to this popularity, the use of AICCs in health and wellbeing settings has received increasing attention (Wester et al., 2024; Doyle et al., 2023; Zhu et al., 2025; Ng and Zhang, 2025; Steenstra and Bickmore, 2025). Evidence suggests that AICCs can help reduce feelings of loneliness, ease emotional distress, enhance mood, and promote self-reflection (Casu et al., 2024; Jung et al., 2023)Skjuve et al. conducted interviews with Replika users and found that users often perceive the chatbot as non-judgmental, which encourages open self-disclosure without fear of social risk. They also found that interactions with the chatbot can be personally rewarding, with users reporting positive effects on their perceived wellbeing (Skjuve et al., 2021).
@@
Despite these benefits, recent work has also raised critical concerns regarding potential harms (Chandra et al., 2025; Zhang et al., 2025; Yoo et al., 2025; Pataranutaporn et al., 2025). Emotional dependence on chatbots can mimic unhealthy attachment patterns (Laestadius et al., 2024; Xie et al., 2023), and users often report disillusionment when chatbots fail to meet expectations (Laestadius et al., 2024). While users often report feelings of social support and wellbeing, this emotional bond may lead to addictive behaviors, where users prioritize interactions with the chatbots over real-life relationships (Marriott and Pitardi, 2024). Notably, although some participants report high perceived social support, this is often situated within contexts of heightened loneliness, particularly among populations already at risk of social isolation (Maples et al., 2024; Liu et al., 2025). A recent study shows that intensive use and frequent self-disclosure are linked to lower wellbeing, particularly among those lacking offline support (Zhang et al., 2025).
@@
While prior research has provided valuable insights, most evaluations of chatbot experiences have relied primarily on interviews and surveys (Dosovitsky and Bunge, 2021; Brandtzaeg et al., 2022; Skjuve et al., 2022; De Freitas et al., 2025b). Although these approaches capture users’ subjective impressions, they may suffer from recall bias and fail to reflect the long-term impacts of chatbot interactions. Recent work has conducted thematic analysis on social media data, offering additional perspectives on real-world experiences (Laestadius et al., 2024; Ta et al., 2020). Our work contributes to this body of work by leveraging longitudinal and large-scale social media data to conduct causal-inference examination of the long-term psychosocial effects of AICCs.
@@ 2.2. Loneliness, Grief, and Mental Health Interventions
Loneliness refers to a distressing feeling due to the perceived absence of certain meaningful relationships that fulfill the individual’s social and emotional needs, rather than from physical isolation (Weiss, 1975). The experience of loneliness is common, particularly in the context of major life changes such as bereavement, where individuals often feel a pronounced sense of social disconnection (Folker et al., 2021; Heinrich and Gullone, 2006). Prior work has linked loneliness to a range of negative wellbeing outcomes, such as sleep disturbances (Griffin et al., 2020), elevated risk of mild cognitive impairment and dementia (Lara et al., 2019), and Alzheimer’s disease (Sundström et al., 2020). Psychologically, loneliness has been found to correlate with increased risk of depression (Erzen and Çikrikci, 2018), anxiety (Sündermann et al., 2014), suicidal ideation (McClelland et al., 2020), and other mental health outcomes (Leigh-Hunt et al., 2017; Moran et al., 2024; Hancock et al., 2022). Among the many contexts where loneliness arises, bereavement is particularly profound. The loss of a loved one not only deepens grief but also increases risks of depression, post-traumatic stress disorder, and other mental and physical health complications (Simon et al., 2014; Erzen and Çikrikci, 2018; Asch et al., 2021; Vedder et al., 2021, 2022).
Amid rising concerns over loneliness and grief, HCI and digital mental health research have turned to digital interventions—ranging from grief-specific tools to broader solutions for social disconnection (Massimi and Baecker, 2010; Doyle et al., 2024; Xygkou et al., 2023; She et al., 2021; Brubaker et al., 2019; Gach and Brubaker, 2021). Prior work has studied mobile applications for self-guided grief support (Doyle et al., 2024; Xygkou et al., 2023; Lei et al., 2025), grief-specific technologies such as digital memorial platforms (She et al., 2021), legacy-building tools, and apps that provide structured guidance through the emotional stages of grief (Baglione et al., 2018, 2017; Doyle et al., 2024).In parallel, digital interventions have emerged to address loneliness (Brubaker et al., 2019). Social networking platforms and purpose-built online communities facilitate peer support and social connectedness for older adults and marginalized groups (Gabarrell-Pascuet et al., 2024), while specialized apps deliver structured activities, psychoeducation, and cognitive-behavioral strategies to help users build social skills and reframe negative thoughts (DuPont et al., 2023; Sharma et al., 2024).
@@ 2.3. Social Media Analysis and Mental Health
Our work is motivated by this growing body of research on social media and mental health experiences. In particular, we examine the impact of engaging with AICCs, such as Replika. Drawing on studies that apply natural language processing and causal inference, our work develops a computational framework to evaluate how dose interactions with AICCs influence users’ emotional expression and psychosocial wellbeing. By analyzing large-scale, longitudinal data, we aim to uncover how these AICCs may function as emotional support and contribute to meaningful changes in users’ mental health trajectories.
@@ 3. Overview of Study Design and Data
The table lists paraphrased Reddit posts from r/replika, illustrating users’ experiences of companionship, emotional support, cost concerns, preferred interaction roles, and perceptions of Replika.
Table 1. Paraphrased Reddit posts in subreddit r/replika
@@
To understand the psychosocial impacts of engaging with AICCs, our study combined large-scale quasi-experimental analyses of public Reddit data with in-depth, semi-structured interviews with active AICC users. This design enabled us to assess longitudinal changes in affective, behavioral, and cognitive dimensions of psychosocial wellbeing, as well as to understand how they perceived and experienced these psychosocial impacts in the context of their everyday lives.
Why Reddit?  Reddit is a widely used social media platform where individuals can pseudonymously participate in topic-based communities, known as “subreddits”. Its pseudonymous design and topic-centered structure are known to support candid self-disclosure, seek social connections, and share life experiences (Kim et al., 2023; De Choudhury and De, 2014; Andalibi et al., 2018; Saha et al., 2019c). Reddit provides a naturalistic and unobtrusive data source for examining real-world behaviors over extended periods at scale, thereby mitigating key limitations of self-report surveys such as recall bias, social desirability, and cross-sectional snapshots (Tourangeau et al., 2000). Prior research has shown that the platform’s longitudinal data enables reliable analysis of behavioral change over time (Saha et al., 2022, 2025b; De Choudhury et al., 2016; De Choudhury and Kıcıman, 2017). Several subreddits focus specifically on AICC use. Among them, r/replika is the largest and includes more than 82,000 members as of December 2025. This community operates as a public space in which users describe their daily interactions with AICCs and reflect on the personal and emotional roles these systems play (Zhang et al., 2024). These characteristics make Reddit an appropriate platform for large-scale and longitudinal examination of how engagement with AICCs relates to changes in psychosocial wellbeing.
We constructed four groups for comparison—1) a Treatment group consisting of users who actively engaged with AICCs (e.g., Replika), 2) a Control-LLM group of generative LLM-based chatbot users who engaged with chatbots (e.g., ChatGPT, Claude, and Gemini), 3) a Control-VA group of general users who actively engaged with AI voice assistants (e.g., Amazon Alexa, Google Assistant), and 4) a baseline Control-NAI group of Reddit users who did not post about AI assistants at all. Drawing on prior literature, we assessed changes in affective, behavioral, and cognitive outcomes before and after engagement using natural language and statistical analyses (Saha et al., 2018; Breckler, 1984; Yuan et al., 2023a). To estimate causal effects, we applied stratified propensity score matching and difference-in-differences regression. Complementing our computational analysis, we conducted semi-structured interviews with 18 active AICC users to understand how they perceived and experienced these psychosocial impacts and the process of developing a relationship with AICCs. Thematic analysis highlighted the diverse ways users experienced support, personal growth, and potential risks in their relationships with AI companions. Finally, we triangulated our quantitative and qualitative findings by contextualizing with literature in psychology and social science (Knapp, 1978). In this section, we describe the data collection process.